OpenStudy (maheshmeghwal9):

If sinx=2cosx Then, what is the value of sin2x?

OpenStudy (anonymous):

x = inv tan (2)

OpenStudy (maheshmeghwal9):

But answer is 4/5.Why?

OpenStudy (dumbcow):

\[\sin(2x) = 2\sin(x) \cos(x)\] replace sin(x) with 2cos(x) \[\sin(2x) = 4\cos^{2} (x)\] solve for cos^2 using triangle  yields cos^2 = 1/5 \[\rightarrow \sin(2x) = 4*\frac{1}{5} = \frac{4}{5}\]